OpenStudy (anonymous):

What fraction is equal to 14/15 – 4/15 ? 6/5 3/5 10/0 2/3

OpenStudy (anonymous):

14/15-4/15=10/15 divide top and bottom by 5: 10/15=2/3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem :)